OpenStudy (anonymous):

how is adding and subtracting rational expressions similar to adding and subtracting simple fractions, and how they are different.

OpenStudy (amistre64):

rational expressions are just glorified fractions ... you can consider a fraction to be a rational expression of constant functions. The operations are exactly the same.

OpenStudy (amistre64):

let a(x) = 2 let b(x) = 3 let c(x) = 5 let d(x) = 1 \[\frac{a(x)}{b(x)}-\frac{c(x)}{d(x)}\] \[\frac{2}{3}-\frac{5}{1}\]